# Quantifying the Cross-sectoral Intersecting Discrepancies within Multiple Groups Using Latent Class Analysis Towards Fairness

### v2

@@ 4.2 EVENS
It is important to note that the EVENS dataset uses a different definition for mixed or multiple ethnic groups compared to Anonymous project. To avoid inconsistencies that could affect the final analysis, we excluded mixed or multiple ethnic groups from the EVENS calculations. Due to the different questions in the EVENS and Anonymous project surveys, we selected three types of cross-sectoral questions from EVENS for analysis: housing, experiences of harassment, and financial situation. These questions were chosen to provide a broad view of participants’ living conditions, social experiences, and economic status during the pandemic.
We applied LCA separately to the data from England and Scotland to uncover patterns and discrepancies within and between these regions. This approach allows us to identify distinct subgroups within the ethnic communities based on their responses to the selected questions, providing deeper insights into the intersectional and cross-sectoral experiences of these groups.
Seeing the discrepancies in the EVENS England data, shown in the left part of Table 2, it is clear that all discrepancy values are relatively small compared to the Anonymous project results. Notably, there are large discrepancies between the Indian and Bangladeshi groups, and between the Indian and Chinese groups. This differs slightly from the conclusions of the Anonymous project’s experiment, likely due to the different types of cross-sectoral questions selected. In the AVG column, the Bangladeshi group has the highest values, indicating they experienced COVID-19 differently. The Census 2021 reported that COVID-19 mortality rates were highest for the Bangladeshi group, for both males and females [7], supporting our findings. The right part of Table 2 describes the discrepancies for the EVENS Scotland data. There are significant discrepancies between the Caribbean and other ethnic groups, likely related to the unique background of the Caribbean group.
Table 2: The matrix of discrepancies for EVENS England and Scotland data.
@@
4.3 Census 2021 (England and Wales)
Table 3: The matrices for the discrepancies of Census 2021 and the deprivation discrepancies between LSOAs across user-defined ME population percentage groups based on Deprivation 2019.
@@ 4.3 Census 2021 (England and Wales)
The proposed approach quantifies the discrepancies between the defined ME population-related groups based on the selected Census 2021 data. The results are shown in the left part of Table 3. It is evident that the discrepancies between LSOAs increase as differences in ME population percentages increase, indicating significant disparities in living conditions for ME individuals across different LSOAs, particularly in terms of energy, housing, and health aspects. Notably, the 0%-20% group shows the largest AVG discrepancy compared to other groups, suggesting that White individuals in those LSOAs experience significantly different living conditions. Since the 0%-20% group constitutes a large portion of the UK (see Appendix B), this finding suggests potential unequal treatment and possible neglect of other LSOAs. Additionally, the 40%-60% group has the smallest AVG discrepancy value, likely due to its intermediate position among the ME groups, sharing characteristics with the 0%-20% and 20%-40% groups, as well as 60%-80% and 80%-100% groups.
@@
Correlation Analysis Furthermore, based on the deprivation indices from Deprivation 2019 [12] and the defined groups, we calculated the percentages of LSOAs in each deprivation-labelled group across various ME population groups. The results are shown in Appendix B, and we treated each row as a feature vector representing one group of LSOAs. We then iteratively calculated the deprivation discrepancies for each pair of rows, with the results presented in the right part of Table 3. We observed similar patterns (color change) to those in the left part of Table 3, which can verify the reliability of our proposed approach.
@@
To statistically verify our proposed approach, we ran Pearson and Spearman row-wise correlation analyses for Census 2021 discrepancies (the left part of Table 3) and Deprivation 2019 discrepancies (the right part of Table 3). The detailed results are shown in Appendix B. All rows exhibit very strong correlations, implying that our approach can draw conclusions very similar to those of experts. Furthermore, we also flattened both matrices to run a one-time correlation analysis. The Pearson correlation coefficient is 0.9797 with a $p$-value of 1.4437e-17, and the Spearman correlation coefficient is 0.9872 with a $p$-value of 7.436e-20. Both $p$-values are far less than 0.001, indicating a strong correlation.
Discrepancy for AI fairness Now, we will show how discrepancies relate to potential AI bias. We selected logistic regression to classify deprivation indices for LSOAs using the Census 2021 data from previous experiments. To simplify the classification task, we redefined the deprivation indices as deprived (indices 1-5, labelled as 0) and not deprived (indices 6-10, labelled as 1). We split the dataset into training and validation sets in an 8:2 ratio. The prediction accuracy on the Census 2021 dataset, with an overall accuracy of 90.35% and a standard deviation (STD) of 4.20 across five group results, is shown in the Census column of Table 4. Meanwhile, the accuracy for each group is displayed on the left in Figure 4. We noticed that the accuracy varies across different groups, with the 0-20% group showing 100% accuracy. In the context of LNOB and AI fairness, we consider this biased and problematic. Additionally, in this research, the STD is considered an important indicator of fairness; smaller STD values imply that the model treats each group more equally. In the following paragraph, we will discuss how the discrepancy relates to AI fairness based on two undersampled datasets.
For the Census 2021 data, we noticed two data imbalance issues likely affecting accuracy. Firstly, the 0-20% group constitutes 82.59% of the LSOA samples and has the largest AVG discrepancy and relatively low accuracy, potentially indicating distinct features compared to other groups. Generally, in machine learning, imbalanced data can negatively impact the majority, leading to issues like overfitting. To address this, we used random undersampling [15], resampling all classes except the minority 80-100% group, which had 125 samples. After undersampling, all groups had an equal number of samples, and we split them into training and test sets with a ratio of 8:2. This method aimed to observe changes in discrepancies, accuracy, and STD to gain insights into bias. The results are shown in the middle of Figure 4 and in the Undersampled Census (ME) column of Table 4.
We found that all AVG discrepancies increased along with the STD (from 4.2 to 4.83), and the prediction accuracy decreased for three groups. This may indicate that our approach efficiently detects data discrepancies that can exacerbate bias issues. Notably, the 40-60% group showed the most significant increase in discrepancy, accompanied by a dramatic decrease in accuracy. We believe that this phenomenon is due not only to the increase in discrepancies but also to the reduced data size, which may not provide enough data for effective model training. When samples from different groups are far apart, the model may require more data to train effectively; otherwise, it may focus on some subareas of the feature space. It is worth noting that while the overall accuracy increases, the accuracy of most groups decreases due to a statistical artifact related to changes in test data size. In this research, our primary focus is on the STD and discrepancies.
Furthermore, another data imbalance issue may be related to deprivation labels. Specifically, in the original Census 2021 data, all groups except the 0-20% group have more data labelled as 0 than 1. Additionally, the ratio of data labelled as 0 to 1 increases with the percentage of the ME population. For example, in the 40-60% group, the number of samples for 0 and 1 labels are 1,014 and 215, respectively, and 537 and 54 for the 60-80% group. We believe this may contribute to bias issues. Therefore, we conducted undersampling on the original Census 2021 data, targeting the labels. After undersampling, the dataset had an equal number of samples for 0 and 1 labels. Then, we split the dataset into training and test sets following the same rule. The results are shown on the right of Figure 4 and in the Undersampled Census (deprivation) column of Table 4.
Overall, we found that all AVG discrepancies decreased, indicating that samples from different groups became closer to each other. This suggests that the model is likely to treat them more similarly in the feature space. On the right of Figure 4, we observed that accuracy across different groups became more similar compared to the other two figures (left and middle), supported by the STD dramatically decreasing to 1.82. In the context of LNOB, the model is now less biased. Meanwhile, the overall accuracy is better than the original one. In general, this set of experiments demonstrated that data discrepancies obtained from our proposed approach can potentially indicate the degree to which the model treats each group equally.
Figure 4: Prediction accuracy for classification tasks on the Census 2021 data (left), undersampled data targeting ME groups (middle), and undersampled data targeting deprivation labels (right).
Table 4: The results of discrepancies and model performance on the original Census dataset and two undersampled Census datasets.
@@
roweven = c, row3 = c, row5 = c, row7 = c, cell12 = c, cell13 = c, cell14 = c, cell15 = c, cell21 = r=2, cell41 = r=5, hline1,9 = -0.08em, hline2 = 2l, hline2 = 3-4, hline2,4 = 5r, hline4 = 1-4, & Census Undersampled Census (ME) Undersampled Census (deprivation)
Model Accuracy $\uparrow$ 90.35 91.20 90.75
STD $\downarrow$ 4.20 4.83 1.82
Discrepancy AVG 0-20% 0.5920 0.6398 0.4966
20-40% 0.3147 0.3242 0.1428
40-60% 0.2414 0.3340 0.1394
60-80% 0.2831 0.3278 0.1584
80-100% 0.3620 0.4340 0.1787
5 Conclusion and Limitations
@@ Appendix A The Anonymous project
As part of the Anonymous project, a multilingual online survey, available in 10 languages, was conducted to investigate the experiences of individuals from minority ethnic groups with digitalised housing, health, and energy services. The survey contains a total of 32 questions. The survey data includes 594 responses from England and 284 responses from Scotland. In terms of respondent selection, researcher carefully determined the required number of participants from each ethnic group in England using a proportional allocation method based on their respective population percentages from the 2021 Census (England & Wales). However, due to the unavailability of Scotland’s 2021 census results during the planning phase, the project aimed to limit the number of respondents from each ME group in Scotland to a maximum of 40. The total number of survey respondents was 878, with 594 participants from England and 284 from Scotland. A detailed breakdown of the respondents’ demographic information, including their ethnicities, is provided in Table 5.
Table 5: The number of respondents from each ME group in both England and Scotland
@@
We show the distribution of responses from seven ethnic groups in England regarding health and digital services based on Anonymous project’s data. We observed a notable discrepancy in the Chinese group, with 30.16% lacking English proficiency and 26.98% struggling to use the online system, while most other ethnic groups reported no concerns.
Figure 5: This figure gives an example of the exploratory data analysis (EDA) for the Anonymous project’s England data focuses on Question 21: “Which of the following concerns do you have about communicating with your general practice (GP) through apps, websites, or other online services?" The bars represent the portion of respondents from each ethnic group selecting each option.
It should be noted that the definitions of ethnicities differ between the Scottish and English censuses. For instance, while individuals in the English census may choose “Pakistani" as a sub-category under the broader category “Asian or Asian British”, in Scotland, this category translates to‘ ‘Pakistani, Pakistani Scottish, or Pakistani British". For consistency, this paper adopts the ethnicity naming conventions used in the English census.
The questions we selected for quantifying discrepancies include: Question 20: Which of the following concerns do you have about communicating with your GP through apps, websites or other online services? Question 27: Do you have any concerns about using an app, website or digital system to carry out energy-related activities? Question 23: Do you have any concerns about using an app, website or other digital service for these housing-related activities?
Figure 6: The example of hyperparameter optimisation process to seek the elbow point.
Regarding the selection of the number of latent classes, we used grid search for hyperparameter optimisation with 10-fold cross-validation. For all experiments, the search space was adjusted based on our empirical observations. In the Anonymous project experiments, we set the range from 2 to 10. For EVENs and Census experiments, we set the range from 2 to 30.

### v3

@@ IV-B EVENS
It is important to note that the EVENS dataset uses a different definition for mixed or multiple ethnic groups compared to the Anonymous project. To avoid inconsistencies that could affect the final analysis, we excluded mixed or multiple ethnic groups from the EVENS calculations. Due to the different questions in the EVENS and Anonymous project surveys, we selected three types of cross-sectoral questions from EVENS for analysis: housing, experiences of harassment, and financial situation. These questions were chosen to provide a broad view of participants’ living conditions, social experiences, and economic status during the pandemic.
We applied LCA separately to the data from England and Scotland to uncover patterns and discrepancies within and between these regions. This approach allows us to identify distinct subgroups within the ethnic communities based on their responses to the selected questions, providing deeper insights into the intersecting cross-sectoral experiences of these groups.
Seeing the discrepancies in the EVENS England data, shown in the left part of Table II, it is clear that all discrepancy values are relatively small compared to the Anonymous project results. Notably, there are large discrepancies between the Indian and Bangladeshi groups, and between the Indian and Chinese groups. This differs slightly from the conclusions of the Anonymous project’s experiment, likely due to the different types of cross-sectoral questions selected. In the AVG column, the Bangladeshi group has the highest values, indicating they experienced COVID-19 differently. The Census 2021 reported that COVID-19 mortality rates were highest for the Bangladeshi group, for both males and females [32], supporting our findings. The right part of Table II describes the discrepancies for the EVENS Scotland data. There are significant discrepancies between the Caribbean and other ethnic groups, likely related to the unique background of the Caribbean group.
TABLE II: The matrix of discrepancies for EVENS England and Scotland data.
@@
IV-C Census 2021 (England and Wales)
TABLE III: The matrices for the discrepancies of Census 2021 and the deprivation discrepancies between LSOAs across user-defined ME population percentage groups based on Deprivation 2019.
@@ IV-C Census 2021 (England and Wales)
The proposed approach quantifies the discrepancies between the defined ME population-related groups based on the selected Census 2021 data. The results are shown in the left part of Table III. It is evident that the discrepancies between LSOAs increase as differences in ME population percentages increase, indicating significant disparities in living conditions for ME individuals across different LSOAs, particularly in terms of energy, housing, and health aspects. Notably, the 0%-20% group shows the largest AVG discrepancy compared to other groups, suggesting that White individuals in those LSOAs experience significantly different living conditions. Since the 0%-20% group constitutes a large portion of the UK (see Appendix B), this finding suggests potential unequal treatment and possible neglect of other LSOAs. Additionally, the 40%-60% group has the smallest AVG discrepancy value, likely due to its intermediate position among the ME groups, sharing characteristics with the 0%-20% and 20%-40% groups, as well as 60%-80% and 80%-100% groups.
@@
Correlation Analysis Furthermore, based on the deprivation indices from Deprivation 2019 [34] and the defined groups, we calculated the percentages of LSOAs in each deprivation-labelled group across various ME population groups. The results are shown in Appendix B, and we treated each row as a feature vector representing one group of LSOAs. We then iteratively calculated the deprivation discrepancies for each pair of rows, with the results presented in the right part of Table III. We observed similar patterns (color change) to those in the left part of Table III, which can verify the reliability of our proposed approach.
@@
To statistically verify our proposed approach, we ran Pearson and Spearman row-wise correlation analyses for Census 2021 discrepancies (the left part of Table III) and Deprivation 2019 discrepancies (the right part of Table III). The detailed results are shown in Appendix B. All rows exhibit very strong correlations, implying that our approach can draw conclusions very similar to those of experts. Furthermore, we also flattened both matrices to run a one-time correlation analysis. The Pearson correlation coefficient is 0.9797 with a $p$-value of 1.4437e-17, and the Spearman correlation coefficient is 0.9872 with a $p$-value of 7.436e-20. Both $p$-values are far less than 0.001, indicating a strong correlation.
Additionally, to further assess the performance of LCA, we conducted experiments using $k$-Means as a replacement for LCA to calculate the discrepancies and performed the same correlation analysis. While the correlation coefficients are still strong, they are lower than those achieved by LCA. This demonstrates the efficiency of LCA and further indirectly supports the usability of the proposed framework.
TABLE IV: The Discrepancies and False Positive Rate.
@@
cells = c, vline2 = -, hline1,8 = -0.08em, hline2,7 = -, & 0-20% 20-40% 40-60% 60-80% 80-100%
0-20% 0.0000 0.5021 0.6775 0.8646 0.9301
20-40% 0.5021 0.0000 0.1272 0.4158 0.5931
40-60% 0.6775 0.1272 0.0000 0.1521 0.3217
60-80% 0.8646 0.4158 0.1521 0.0000 0.0450
80-100% 0.9301 0.5931 0.3217 0.0450 0.0000
FPR 0.1225 0.0710 0.0421 0.0090 0
TABLE V: Discrepancy Values for Five Groups with Three Sampling Ratios.
{tblr}
cells = c, cell11 = r=2, cell31 = r=5, cell81 = r=5, vline8,13 = -0.05em, hline1,13 = -0.08em, hline2 = 2-170.03em, hline3,8 = -0.05em, Model & Sampling Ratio 80% 90% 100%
Ethnic Group 1 2 3 4 5 1 2 3 4 5 1 2 3 4 5
LR 1 0.0000 0.1407 0.3530 0.5661 0.8130 0.0000 0.1424 0.3558 0.5698 0.8141 0.0000 0.1429 0.3542 0.5675 0.8120
2 0.1407 0.0000 0.0955 0.3095 0.6872 0.1424 0.0000 0.0979 0.3141 0.6872 0.1429 0.0000 0.0966 0.3115 0.6853
3 0.3530 0.0955 0.0000 0.1061 0.5593 0.3558 0.0979 0.0000 0.1051 0.5553 0.3542 0.0966 0.0000 0.1043 0.5541
4 0.5661 0.3095 0.1061 0.0000 0.4005 0.5698 0.3141 0.1051 0.0000 0.3983 0.5675 0.3115 0.1043 0.0000 0.3971
5 0.8130 0.6872 0.5593 0.4005 0.0000 0.8141 0.6872 0.5553 0.3983 0.0000 0.8120 0.6853 0.5541 0.3971 0.0000
MLP 1 0.0000 0.1421 0.3520 0.5663 0.8111 0.0000 0.1426 0.3535 0.5681 0.8118 0.0000 0.1425 0.3533 0.5679 0.8119
2 0.1421 0.0000 0.0960 0.3120 0.6853 0.1426 0.0000 0.0964 0.3135 0.6856 0.1425 0.0000 0.0967 0.3139 0.6865
3 0.3520 0.0960 0.0000 0.1059 0.5546 0.3535 0.0964 0.0000 0.1068 0.5548 0.3533 0.0967 0.0000 0.1066 0.5554
4 0.5663 0.3120 0.1059 0.0000 0.3955 0.5681 0.3135 0.1068 0.0000 0.3958 0.5679 0.3139 0.1066 0.0000 0.3967
5 0.8111 0.6853 0.5546 0.3955 0.0000 0.8118 0.6856 0.5548 0.3958 0.0000 0.8119 0.6865 0.5554 0.3967 0.0000
TABLE VI: False Positive Rates for Five Groups with Three Sampling Ratios.
{tblr}
cells = c, vline5 = -0.05em, hline1,8 = -0.08em, hline2-3 = -0.05em, Model & Logistic Regression Multilayer Perceptron
Sampling Ratio 80% 90% 100% 80% 90% 100%
1 0.1285 ± 0.0161 0.1367 ± 0.0135 0.1383 ± 0.0144 0.1159 ± 0.0513 0.1464 ± 0.0525 0.1384 ± 0.0378
2 0.118 ± 0.018 0.1124 ± 0.0147 0.1127 ± 0.012 0.0872 ± 0.0304 0.1234 ± 0.0318 0.1067 ± 0.0306
3 0.1259 ± 0.0131 0.1272 ± 0.014 0.1277 ± 0.0177 0.1 ± 0.026 0.1341 ± 0.043 0.1173 ± 0.0268
4 0.1232 ± 0.0175 0.1298 ± 0.0113 0.1231 ± 0.0083 0.0948 ± 0.0257 0.1327 ± 0.0342 0.1046 ± 0.0275
5 0.0603 ± 0.0072 0.0624 ± 0.0058 0.0623 ± 0.0057 0.0469 ± 0.0193 0.0637 ± 0.0199 0.0533 ± 0.0175
V Discrepancy for AI fairness
As we discussed earlier, our proposed approach can be used to quantify the discrepancies in the distribution of user (sample) profiles across multiple groups. We consider that these discrepancies can negatively impact the fairness of machine learning methods. Intuitively, the training process of machine learning models may struggle to extract patterns equally from two parts of a dataset with significant discrepancies. We expect that the discrepancy can serve as a data exploratory metric to alert AI users to the risk of fairness issues and support fairness analysis in AI.
V-A Predefined Range Groups (Fixed Intervals)
Now, we will show how discrepancies relate to potential AI bias. We selected logistic regression (LR) to classify deprivation indices for LSOAs using the Census 2021 data from previous experiments. To simplify the classification task, we redefined the deprivation indices as deprived (indices 1-5, labelled as 0) and not deprived (indices 6-10, labelled as 1). We randomly split the dataset into training and validation sets in an 8:2 ratio, and the experiments are repeated for 10 times.
The bias is measured using the False Positive Rate (FPR). In this study, we argue that FPR deserves greater attention, as it quantifies the extent to which deprived areas are incorrectly predicted as not deprived. Such misclassification could potentially exacerbate deprivation. For instance, groups with higher FPR may not receive the deserved attention they need from the government when it comes to resource allocation decisions.
The experimental results indicate that the two groups with the largest discrepancy values exhibit the greatest difference in FPR. As shown in Table IV, the discrepancy between the 0-20% group and the 80-100% group is the largest with the value 0.9301. At the same time, the former group has the smallest FPR of 0, while the latter group has the largest FPR of 0.1225. This study reveals that the machine learning model can treat areas predominantly populated by non-ethnic minorities unfairly. Moreover, significant discrepancies between groups within a dataset warrant attention, as these disparities may lead to differential treatment by machine learning models.
V-B Equal-Size Groups (Quantile-Based)
We observed that the grouping method employed in the aforementioned experiments presents a data imbalance issue, as the 0-20% group constitutes 82.59% of the LSOA samples. This imbalance may undermine the robustness of our findings, given the significant decrease in the number of samples from the 0-20% group to the 80-100% group. Consequently, the 80-100% group may achieve a 0 FPR with a small number of test samples, despite having a limited amount of training data.
To address this issue, we propose an alternative grouping method. First, we calculate the ME population percentages for all LSOAs, then sort and divide them into five groups, ensuring each group contains an equal number of samples. The ME population percentage ranges for these groups are as follows: 0-0.96%, 0.96-2.33%, 2.33-6.16%, 6.16-17.34%, and 17.34-95.02%. These groups are labeled as 1, 2, 3, 4, and 5, respectively. It is worth noting that the average ME population percentage in the UK is 18%, indicating that only the last group can be considered representative of the ME population.
To further explore the correlations between discrepancy and AI fairness, we conducted experiments using two models: LR and a Multilayer Perceptron (MLP). Additionally, we applied two sampling ratios (90% and 80%) to randomly generate two new datasets. From our perspective, the sampled datasets should preserve the patterns discussed in the previous section, as random sampling does not substantially alter the overall distribution. All the results presented in Table V and Table VI are derived from experiments repeated 10 times.
In Table V, we observed that the discrepancies between Groups 1, 2, 3, 4, and 5 have increased compared to the values presented in Table IV. We believe this is due to all ME LSOAs being concentrated in group 5 under the new grouping method, which explains this observation. In other words, the ME group and non-ME groups have differences in their user profiles. Meanwhile, the largest discrepancy remains between group 1 (predominantly non-ME) and Group 5 (predominantly ME) across the two sampled datasets (70% and 80%) and the original dataset (100%). In Table V, “LR” and “MLP” indicate that the discrepancy calculations were performed on independently sampling datasets. The datasets are used to observe the correlations between discrepancy and AI bias.
In Table VI, the first group (predominantly non-ME), approximately equivalent to the 0-20% group in Table IV, still exhibits the largest FPR across both models and all three datasets. Meanwhile, Group 5 continues to have the smallest FPR. Thus, the results align with the findings presented in Section V-A; the two groups with large discrepancy values may be treated differently by AI.
Additionally, compared to the results shown in Table IV, we observe that the FPR values for Groups 2$\sim$4 increase along with the increase of discrepancy values between Groups 2$\sim$4 and Group 5. Meanwhile, the overall discrepancies within Groups 2$\sim$4 are smaller than the discrepancies between Group 5 and the other groups (Groups 1$\sim$4). This indicates that the features of the data for Groups 2$\sim$4 are relatively similar, and models treat them similarly. As shown in Table IV, Groups 1$\sim$4 received relatively similar FPRs, while the FPR for Group 5 is significantly smaller. In summary, we believe our proposed method effectively quantifies the discrepancies between different groups. Additionally, these values are important for informing AI users and highlight potential risks of AI unfairness.
VI Conclusion and Limitations
@@ Appendix A The Anonymous project
As part of the Anonymous project, a multilingual online survey, available in 10 languages, was conducted to investigate the experiences of individuals from minority ethnic groups with digitalised housing, health, and energy services. The survey contains a total of 32 questions. The survey data includes 594 responses from England and 284 responses from Scotland. In terms of respondent selection, researcher carefully determined the required number of participants from each ethnic group in England using a proportional allocation method based on their respective population percentages from the 2021 Census (England & Wales). However, due to the unavailability of Scotland’s 2021 census results during the planning phase, the project aimed to limit the number of respondents from each ME group in Scotland to a maximum of 40. The total number of survey respondents was 878. A detailed breakdown of the respondents’ demographic information, including their ethnicities, is provided in Table A.1.
TABLE A.1: The number of respondents from each ME group in both England and Scotland
@@
We show the distribution of responses from seven ethnic groups in England regarding health and digital services based on Anonymous project’s data. We observed a notable discrepancy in the Chinese group, with 30.16% lacking English proficiency and 26.98% struggling to use the online system, while most other ethnic groups reported no concerns, as shown in Fig. A.1.
Figure A.1: This figure gives an example of the exploratory data analysis (EDA) for the Anonymous project’s England data focuses on Question 21: “Which of the following concerns do you have about communicating with your general practice (GP) through apps, websites, or other online services?” The bars represent the portion of respondents from each ethnic group selecting each option.
Figure A.2: The example of hyperparameter optimisation process to seek the elbow point.
It should be noted that the definitions of ethnicities differ between Scotland’s census and the England & Wales census. For instance, in the England and Wales census, individuals may select “Pakistani” as a sub-category under the broader “Asian or Asian British” category, whereas in Scotland census, this category is listed as “Pakistani, Pakistani Scottish, or Pakistani British.” For consistency, this paper adopts the ethnicity naming conventions used in the England and Wales census.
The questions selected to quantify discrepancies in this study include:
  • Question 20: Which of the following concerns do you have about communicating with your GP through apps, websites or other online services?
  • Question 23: Do you have any concerns about using an app, website or other digital service for these housing-related activities?
  • Question 27: Do you have any concerns about using an app, website or digital system to carry out energy-related activities?
Regarding the selection of the number of latent classes, we used grid search for hyperparameter optimisation with 10-fold cross-validation. The search space for all experiments was adjusted based on empirical observations. In the Anonymous project experiments, we set the range from 2 to 10, while for EVENs and Census experiments, it was set from 2 to 30. One example result of hyperparameter optimisation is shown in Fig. A.2.